GreenHab Report – <date – November 10th>

Report title: GreenHab Report
Crew # 322
Position: GreenHab Officer
Report prepared by: Leszek Orzechowski
Date: 10-11-2025
Sol: 1

Environmental control (fan & heater): GH Heater: 60 F, GH Fan max 80 F, GH Fan min 75
Average temperatures (last 24h): 80 F
Maximum temperature (last 24h): 112 F
Minimum temperature (last 24h): 60F
Hours of supplemental light: : Suplemental lights are turned off
Daily water usage for crops: 20 liters
Daily water usage for research and/or other purposes:
Time(s) of watering for crops: 10 am 6:30 PM
Changes to crops: none to report
Soil Moisture Level: Changed from Moist to Wet
Microgreen Growth Experiment: We will start microgreens tomorrow
Space Nutrition Task: part of the SubTask – 1 Food inventory for the mission complete – next step is to add nutritional values to each food item. Seeds in the green hab were also reviewed for the Space Nutrition Task
Harvest: no harvest to report

Support/supplies needed: would appreciate some instructions for new growbox.

Operations Report – <date – November 10th>

Report title: Operations Report
Crew #: 322
Position: Engineer
Report prepared by: Alessandro Balossino
Date: 10-11-2025
Sol:1

Non-nominal systems: None
Notes on non-nominal systems: None

ROVERS

Spirit rover used:
Hours: 287h
Beginning charge: 100%
Ending charge: 72%
Currently charging: Yes

Opportunity rover used:
Hours: 230.5h
Beginning charge: 100%
Ending charge: 67%
Currently charging: Yes

Curiosity rover used:
Hours: 328.8h
Beginning charge: 100%
Ending charge: 58%
Currently charging: Yes

Perseverance rover used:
Hours: 367.1h
Beginning charge: 100%
Ending charge: 90%
Currently charging: Yes

General notes on rovers: Everything nominal

Summary of Hab operations:
Water Use (see notes): 70 gallons
Main tank level (remaining gallons, see notes): 364
Main water tank pipe heater (ON or OFF): ON
Main water tank heater (ON or OFF): ON
Toilet tank emptied (NO or YES): NO

Summary of internet: all nominal
Summary of suits and radios: all nominal.

Summary of GreenHab operations: (use information from the GreenHab report)
Water use: <1 gallon (160 gallons – 99,8%)
Heater (ON or OFF): ON
Supplemental light (hours of operation): 0
Harvest (name, weight in grams): 0

Summary of ScienceDome operations: N/A
Dual split (Heat or AC, On or Off or Automatic): Automatic

Summary of RAM operations: None
Summary of any observatory issues: None
Summary of health and safety issues: None

Questions, concerns and requests to Mission Support: None

Sol Summary – <date – November 10th>

Report title: Sol Summary Report
Crew #: 322
Position: Co-Commander A
Report prepared by: Malica Schmidt
Date: 10-11-2025
Sol: 1

Summary Title: Sol Summary Report
Mission Status: Nominal
Sol Activity Summary: 3 EVAs

Mission Task Updates:
– Space Nutrition Task: Inventory of food items completed
– VR Training: Not started
– Mission Vlog: Initial planning completed
– Ultrasound: All AAs finished their training and exam

Look Ahead Plan:
– Sol 2: 2 EVAs, ultrasound phantom familiarization and training session for all AAs, ultrasound tele practice, start of VR Training.

Anomalies in work:
General:
Digital Compass did not work.
EVA 1:
– fNIRS scanner did not connect, restarting did not solve the issue.
– fNIRs (8219) did not work. During EVA, one fNIRs fell off during the rover drive to the site, one fNIRs fell off on the way back.
– One pair of Tobii glasses did not work, setup nominal but one was switched to ethernet. This data was lost after finishing the EVA due to – connection error.
– PLSS ventilation failed during the drive.
– Visor fell off.

EVA 2:
– fNIRS (8219) did not pair with laptop and we were not able to start the test.
– Tobii glasses (6881) setup was nominal but disconnected during EVA. Unable to establish connection after finishing EVA, leading to loss of data.
– PLSS ventilation failed during pre-EVA.

EVA 3:
– fNIRS scanner did not connect, restarting solved the issue.
– fNIRS (8219) connected to laptop. Device experienced an electrical issue. Error with Bluetooth antenna but left probe (2) was working.
– Tobii glasses connected using ethernet cable but ran out of power. EV1 pair of glasses changed batteries and captured data. Second did not change and loss of data.

Weather: Sunny with few clouds, maximum 64°F, minimum 30°F
Crew Physical Status: High spirited, stressed for life outside, tired, things started slow but seem to come together
EVA: 3
Reports to be filled: Sol Summary, Journalist Report, GreenHab Report, Operations Report, HSO Report
Support Requested: N/A

Sol Summary – <date – November 9th>

Report title: Sol Summary Report
Crew #: 322
Position: Co-Commander A
Report prepared by: Malica Schmidt
Date: 09-11-2025
Sol: 0

Summary Title: Sol Summary Report
Mission Status: Pre-mission
Sol Activity Summary: Training day
Mission Task Updates:
• 3 Test EVAs for all analog astronauts including testing of tracking glasses, fNIRS,
• Rover test drive for all analog astronauts
• Habitat tour and moving in
Look Ahead Plan: Begin of the mission sol 1 includes 3 EVAs, Ultrasound Training and Exam, Space Nutrition Task, VR Training

Anomalies in work: N/A
Weather: Sunny with no clouds, maximum 64°F, minimum 30°F
Crew Physical Status: High spirited, excited but tired
EVA: 3
Reports to be filed: 6
Support Requested: MCC and Mission support were on-site throughout the day

Mission Plan – <Date – November 9th>

Report title: Mission Plan
Crew #: 322
Position: Experimental Team
Report prepared by: Erin Richardson

Mission Plan:

Our mission objective is to conduct a full-scale Martian analog designed to study decision making under conflicting information in isolated, confined, and extreme environments. This week, we have a diverse, highly skilled, international crew of 6 staying in the habitat. During their stay, the crew will conduct a total of 12 EVAs. The crew will conduct at most 3 EVAs per day, where no crew member ever participates in more than one EVA per day. Throughout the week, the crew members will visit different EVA sites and complete missions with unique objectives, tasks, and tools. Beyond the EVAs, the crew will complete an ultrasound cardiac imaging task, train on Mars-relevant tasks in virtual reality, and conduct a study on space nutrition and supplementation via the GreenHab. We are looking forward to a busy, exciting, and productive week ahead!

Journalist Report – <date – November 9th>

Report title: Journalist Report
Crew #: 322
Position: Scientist
Report prepared by: Anastasia G. Yanchilina
Date: (09-11-2025)
Sol: 0

Journalist Report: Sol 0, the beginning of human search for biosignatures of Mars.

Crew 322 started their first mission on sol 0 at Mars. This largely included training and learning how to use all of the equipment and doing a sample EVA. The EVA was to locate a specific location where the rover was located that found the location of potential biosignatures, that were described as gypsum. The crew did the EVA’s in pairs of two and three total. One of the analogue astronauts had a topographic map and the other had a compass and a tripod (with the tripod noting the signal of the location of the rover). Two of the crews were able to locate the rover and one of the groups didn’t due to issues with communication.

Following the sample training EVA, the group embarked to do training on the rovers they were to be using. Everyone learned very fast and did a fast route out and back for practice.

Operations Report – <date – November 9th>

Report title: Operations Report
Crew #: 322
Position: Engineer
Report prepared by: Alessandro Balossino
Date: 09-11-2025
Sol:0

Non-nominal systems: None
Notes on non-nominal systems: N/A

ROVERS
Spirit rover used:
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Opportunity rover used:
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Curiosity rover used:
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Perseverance rover used:
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

General notes on rovers: used all rovers for a short drive as part of the crew’s training. All nominal.

Summary of Hab operations:
Water Use (see notes): 87.5%*
Main tank level (remaining gallons, see notes): 364*
Main water tank pipe heater (ON or OFF): ON
Main water tank heater (ON or OFF): ON
Toilet tank emptied (NO or YES): NO

*The actual level is higher, since the tank is filled to the top.

Summary of internet: all nominal
Summary of suits and radios: all nominal

Summary of GreenHab operations: (use information from the GreenHab report)
Water use: 0 gallons (160 gallons – 100%)
Heater (ON or OFF): ON
Supplemental light (hours of operation): 0
Harvest (name, weight in grams): 0

Summary of ScienceDome operations: N/A
Dual split (Heat or AC, On or Off or Automatic): N/A

Summary of RAM operations: None
Summary of any observatory issues: None
Summary of health and safety issues: None

Questions, concerns and requests to Mission Support: None

Supplemental Operations Report 08Nov2025

[title Supplemental Operations Report – November 8th]

Date: 11/8/2025
Name of person filing report: Anderson Wilder
Reason for Report: Routine
Non-Nominal Systems: Crew car. Power system battery, inverters. HAB outer shell,

Power system:
1) 200 hr maintenance (including oil, oil filter, and air filter change) expected on 11/12.
2) Current generator hours – 168.

Propane Readings:
Station Tank: 85%
Director Tank: 72%
Intern Tank: 72%
Generator Tank: 38%
Propane refill scheduled for Monday 11/10

Water:
Hab Static Tank – 550 gallons (above safety level 130 gallons)
GreenHab – 200 gallons (above safety level 60 gallons)
Outpost tank – 450 gallons
Septic Tanks emptied: Yes on 10/20/2025

Rovers:
Sojourner rover used: No
Hours: 221.0
Beginning Charge: 100 %
Ending Charge: _ %
Currently Charging: Yes

Cars:
Hab Car used and why, where: To Hanksville for supplies
Crew Car used and why, where: To Hanksville for supplies, Grand Junction for crew pick up.
General notes and comments: Crew car driver’s side front ball joint is bad

Summary of Internet: Nominal.
EVA suits and radios:
Suits: Nominal.
Comms: Nominal.
T-Echo EVA-link: Nominal.

Campus:Nominal.
Summary of Hab Operations: Nominal.
Summary of GreenHab Operations: Nominal.
Summary of SciDome Operations: Nominal.
Summary of Observatories Operations: Nominal.
Summary of RAM Operations: Cleaned. Nominal.
Summary of Outpost Operations: Nominal.
Summary of Health and Safety Issues: Nominal.

Copyright © The Mars Society. All rights reserved. | Main Site